OpenStudy (anonymous):

Solve tghe euation using cramers law 2x-2y-3z=9 3x+z=10 x+y=0

OpenStudy (anonymous):

Solve by Substitution : // Solve equation [3] for the variable y [3] y = -x // Plug this in for variable "y" in equation [1] [1] 2x - 2•(-x ) - 3z = 9 [1] 4x - 3z = 9 // Solve equation [2] for the variable z [2] z = -3x + 10 // Plug this in for variable "z" in equation [1] [1] 4x - 3•(-3x+10) = 9 [1] 13x = 39 // Solve equation [1] for the variable x [1] 13x = 39 [1] x = 3 // By now we know this much : x = 3 y = -x z = -3x+10 // Use the x-value to solve for y y = -(3) = -3 // Use the x-value to solve for z z = -3(3)+10 = 1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@enano64fs Np. :) Just hit the Best response next to me, It helps me out alot :D

OpenStudy (anonymous):

@Loser66 lol what do you mean? I already have 2 medals here